OpenStudy (anonymous):

5s-25/s^2-25 need help

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{5s-25 }{s^2-25} }\) Factor the numerator and the denominator. For the numerator I will tell you that both terms have a "5" in common. For the denominator, use the rule: \(\color{black}{\rm g^2-f^2=(g-f)(g+f)}\)

OpenStudy (anonymous):

5/s+5 would that be my answer

jagr2713 (jagr2713):

Yes thats the answer for the top :D How about the bottom

OpenStudy (nincompoop):

I do not know if that would be your answer, because you didn't even show your step solution

jagr2713 (jagr2713):

Wait what are the instructions ?

OpenStudy (anonymous):

write each rational expression in the lowest terms sorry

OpenStudy (solomonzelman):

you did get the correct answer, I believe :)

OpenStudy (anonymous):

thank you so much . really needed the help

OpenStudy (solomonzelman):

yw

jagr2713 (jagr2713):

Yep